OpenStudy (anonymous):

Estimate the number of pennies you will able to collect in a cylindrical piggy-bank that has a diameter of 3 inches and a height of 6 inches. Help pleasee :(

OpenStudy (anonymous):

pi * 1.5^2 * 6 = 42.4115008 cubic inches VOLUME of penny = 3.14159 * 9.5 * 9.5 * 1.27 = .360 cubic centimeters = 0.0219685479 cubic inches 42.4115008 / 0.0219685479 = 1 930.55549 pennies Check me if I'm wrong

OpenStudy (anonymous):

Aha, so we needed to know the height & radius of a penny

OpenStudy (anonymous):

We did an estimation problem at the beginning of class which was 2 months ago, we didnt use any measurements etc, so i didnt think we would for this one. We didnt do any other estimation problems ever since the beginning of class